# Adaptive path planning for efficient object search by UAVs in agricultural fields

### v1

@@ Abstract
This paper presents an adaptive path planner for object search in agricultural fields using UAVs. The path planner uses a high-altitude coverage flight path and plans additional low-altitude inspections when the detection network is uncertain. The path planner was evaluated in an offline simulation environment containing real-world images. We trained a YOLOv8 detection network to detect artificial plants placed in grass fields to showcase the potential of our path planner. We evaluated the effect of different detection certainty measures, optimized the path planning parameters, investigated the effects of localization errors and different numbers of objects in the field. The YOLOv8 detection confidence worked best to differentiate between true and false positive detections and was therefore used in the adaptive planner. The optimal parameters of the path planner depended on the distribution of objects in the field, when the objects were uniformly distributed, more low-altitude inspections were needed compared to a non-uniform distribution of objects, resulting in a longer path length. The adaptive planner proved to be robust against localization uncertainty. When increasing the number of objects, the flight path length increased, especially when the objects were uniformly distributed. When the objects were non-uniformly distributed, the adaptive path planner yielded a shorter path than a low-altitude coverage path, even with high number of objects. Overall, the presented adaptive path planner allowed to find non-uniformly distributed objects in a field faster than a coverage path planner and resulted in a compatible detection accuracy. The path planner is made available at https://github.com/wur-abe/uav_adaptive_planner.
@@ 1 Introduction
Over the last years, the use of Unmanned Aerial Vehicles (UAVs) in agriculture has grown considerably (Rejeb et al., 2022). Applications are ranging from yield estimation (Yuan et al., 2024) to weed (Murad et al., 2023) and disease detection (Shahi et al., 2023; Kouadio et al., 2023). These applications require objects, such as weeds and diseases, to be detected in a large field. Typically, the UAV follows a pre-defined low-altitude coverage flight path at fixed altitude (Cabreira et al., 2019). These paths are easy to use and cover the whole field with an equal spatial resolution. However, for an object detection task, it is not needed to cover the whole field with an equal spatial resolution. Flying at a higher altitude will increase the field-of-view of the UAV and thereby decrease the total flight path length and associated flight time, but decreases the spatial resolution and thereby the detection accuracy. Only decreasing the altitude of the UAV when the detection is uncertain can decrease the flight time while achieving a similar detection as a low-altitude coverage path. To optimize the battery capacity of UAVs, it is important to use a method that minimizes the flight path length while maintaining a good detection accuracy. Especially when using UAVs for autonomous and frequent monitoring of fields, it is important to limit the flight time to increase the area that can be monitored using a single UAV.
@@
Adaptive path planning, sometimes called informative path planning, is a path planning approach that aims to plan a trajectory that maximizes gathering of task-relevant information while keeping within specific resource limits, such as battery capacity (Popović et al., 2024). This can be done by making the path planning reactive to real-time data, making sequential decisions about the UAV’s flight path based on newly gathered information (Popović et al., 2024). Adaptive path planning with UAV’s is used, for instance, to detect anomalies in ocean topography (Blanchard and Sapsis, 2022), detection of weeds in agricultural fields (Popović et al., 2017) and multi-resolution semantic segmentation (Stache et al., 2023). Popović et al. (2017) showed that an adaptive path planner yielded a 50% lower map uncertainty (measured using entropy) in the same amount of time for active weed classification (simulated using printed markers) compared to a low-altitude coverage flight path. Stache et al. (2023) proposed a path planner for semantic segmentation that estimates the relation between altitude and segmentation accuracy. It performs close inspections at lower altitudes only when required and thereby minimizing the flight time. The path planner was demonstrated by segmenting soil, plants and weeds by artificially flying over the WeedMap dataset and elements like asphalt, buildings, water etc. by artificially flying over satellite images. They show a high segmentation accuracy while minimizing the flight time. van Essen et al. (2024) presents a learning based adaptive path planner that uses a Reinforcement Learning to localize objects of interest in a field. The learned policy uses uncertain prior knowledge of the field together with object detections in the current field-of-view to control a UAV. In a simulated environment, they show that learned path outperforms a coverage planner by localizing 80% of the objects in less than half of the flight time.
@@
When increasing the altitude of the UAV, the field-of-view increases and thereby the spatial resolution of the image decreases. Due to a smaller number of pixels describing an object, the number of false positive and false negative detections will increase and thereby the accuracy decreases. An adaptive path planner can exploit the trade-off between altitude and accuracy, by only capturing low-altitude images when needed. For example, at a high altitude, the detection network may detect an object that appears to be a weed. Based on the uncertainty associated with the detection, a decision can be made to accept the detection as reliable or gather a new image at lower altitude.
@@
A detection network typically outputs a confidence belonging to a detection as measure of uncertainty. When the detection confidence of an object is low, an adaptive path planner can, for example, plan a low-altitude inspection of this object. However, neural networks tend to over- or under-estimate the confidence (Gawlikowski et al., 2023), leading to a rejection of correct detections and an inclusion of wrong detections. Several methods have been applied in literature to provide a better uncertainty estimation. Applying Monte-Carlo Dropout (MCD) during inference is often applied due to its low implementation effort (Gawlikowski et al., 2023). Myojin et al. (2019) showed a 1-2% increase in accuracy when using uncertainty measures estimated using MCD for detecting lunar craters on images using an adapted version of YOLOv3. In agriculture, MCD has been applied for uncertainty estimation in active learning for broccoli detection (Blok et al., 2022), fish species detection (Sokolova et al., 2024), obstacle detection in solar farms (Rodriguez-Vazquez et al., 2024), and assessing uncertainty in chicken plumage condition (Lamping et al., 2023). They illustrate that MCD can be used to estimate uncertainty in neural networks.
@@
The objective of this paper is to develop and evaluate an adaptive path planner that finds objects in agricultural fields faster than a low-altitude coverage path planner while maintaining a similar detection performance. This adaptive path planning method follows a high-altitude coverage flight path and uses the detection certainty to perform low-altitude inspections when the detection certainty was low. We evaluated the effect of different certainty estimation methods calculated using MCD and optimized the flight altitude and certainty thresholds of the path planner. Most existing approaches assume a perfect localization of the UAV, which limits the practical usability (Popović et al., 2024). Therefore we investigated the effect of errors in the UAV’s localization. The number of objects to be detected in the field influences the efficiency of the adaptive path planner, therefore we also investigated this relation. The code for the path planner is made available at https://github.com/wur-abe/uav_adaptive_planner.
@@ 2.1 Adaptive path planner
The concept of the proposed adaptive path planner is visualized in Figure 1. First, the flight area to be covered by the UAV needs to be manually selected. Based on the flight area, a high-altitude coverage path is planned (section 2.1.1) and executed. At each waypoint, we take an image from top-down perspective (nadir) and a detection network is used to detect the objects in real-time (section 2.1.2). The detected objects are then georeferenced (section 2.1.3) and mapped (section 2.1.4). After completing the high-altitude coverage path, an inspection path is planned with waypoints at a lower altitude at the locations of the object detections that were uncertain (section 2.1.5). To limit the number of waypoints for low-altitude inspection, the waypoints are filtered to remove overlapping waypoints. The inspection flight path is executed in a similar way as the high-altitude coverage path. During the inspection flight, the mapped objects are updated and previously unseen objects are added when needed. Finally, all object locations in the map (both from the high- and low-altitude flight) with a certainty larger than $c_{\textrm{eval}}$ are used for evaluation.
@@ 2.1.1 Planning of high-altitude coverage path
The first step of the proposed method is to manually select a flight area by drawing a polygon on a satellite image indicating the boundaries of the field. Within this area, a coverage flight path is planned using Fields2Cover (Mier et al., 2023) at an altitude of $h_{\textrm{cov}}$. The field-of-view, $FOV(h)$ (m) at altitude $h$ is calculated by:
$FOV(h)=\frac{\textbf{c}\odot h}{f},$ (1)
where $\odot$ denotes an element-wise multiplication, $\textbf{c}=[c_{w},c_{h}]^{T}$ the sensor width and height in mm, and $f$ the camera focal length in mm. However, when the localization of the UAV is imperfect, taking the calculated field-of-view as row width for the coverage path planner would then result in missing a part of the field. Therefore, to reduce the impact of localization errors, we set the row width to 90% of the field-of-view by adding an overlap of 5% of the field-of-view to both sides. Likewise, the distance between two waypoints on the flight path is set to 90% of the field-of-view. Figure 2 shows an example of the coverage flight paths for 12m, 24m, 36m and 48m altitude.
Figure 2: Example of the coverage flight path for 12m (a), 24m (b), 36m (c) and 48m(d) altitude.
@@ 2.1.2 Detection
At each waypoint, an image is taken from top-down perspective. In this image, objects are detected using YOLOv8. (Jocher et al., 2023). YOLOv8 is the eight version of the YOLO (You Only Look Once) family of fully convolutional neural networks for object detection, which predicts both the bounding box coordinates as well as the class probabilities in a single pass through the network. The nano version of YOLOv8 is used, which was selected to have low hardware requirements, and allows for detection on-board of a UAV. During inference, an image input size of 2048x2048 pixels is used, and the minimum confidence is set to $c_{\textrm{reject}}$.
@@ 2.1.3 Georeferencing
Georefencing object detections is done in real-time using the UAV’s position and heading. The first step is to calculate the field-of-view, $FOV(h)$, at the current altitude of the UAV using Equation 1. The pixel coordinates of the detected object $o$ in the image $[u_{o},v_{o}]^{T}$ (px) can be transformed in the local camera frame, $[x^{\textrm{local}}_{o},y^{\textrm{local}}_{o}]^{T}$ (m), by:
$\begin{bmatrix}x^{\textrm{local}}_{o}\\ y^{\textrm{local}}_{o}\end{bmatrix}=\Biggl{(}\begin{bmatrix}u_{o}\\ v_{o}\end{bmatrix}\odot\begin{bmatrix}\frac{1}{I_{w}}\\ \frac{1}{I_{h}}\end{bmatrix}-\frac{1}{2}\Biggr{)}\odot FOV(h),$ (2)
@@
where $\psi$ is the counter-clockwise heading of the drone relative with respect to the north and $[x^{\textrm{utm}}_{\textrm{uav}},y^{\textrm{utm}}_{\textrm{uav}}]^{T}$ the UAV’s UTM coordinates. Since UTM coordinates uses the North-up orientation, we take the negative of $y^{\textrm{local}}_{o}$.
@@ 2.1.4 Mapping
The GNNS location of each detection is stored in a map during operation. Because of overlap between waypoints and flying at different altitudes, objects may be observed multiple times. To merge these observations, all detections with an Euclidean distance of less than $d_{\textrm{dist}}$ are considered the same object. The default value of $d_{\textrm{dist}}=0.35$m was selected based on the minimum distance between the objects in the field.
@@ 2.1.5 Planning of low-altitude inspection flight path
After executing the high-altitude coverage flight path, an inspection flight path is planned to observe objects in the map with a low certainty (see section 2.2) in more spatial detail at a lower altitude. For all the objects stored in the map, we make a decision $\mathcal{D}_{i}$ for each object $i$, to accept, reject or have a closer inspection of the object using:
@@
where $c_{i}$ is the certainty corresponding to the object (see section 2.2), $c_{\textrm{accept}}$ the certainty threshold to accept an object without further inspection and $c_{\textrm{reject}}$ the certainty threshold to reject and ignore an object. An inspection flight path is planned at an altitude of $h_{\textrm{inspect}}$ to visit all objects were $\mathcal{D}_{i}$ equals ’inspect’. The resulting list of waypoints is ordered to get the shortest route using iterative stochastic local search by minimizing the total distance (Mulvad, 2022). For every waypoint, the heading is set towards the next waypoint. To shorten the created path, the waypoints are iteratively filtered by removing next waypoints when the corresponding detection is expected to be visible in the field-of-view of the current waypoint. Figure 3 shows an example of an inspection flight path.
@@ 2.2 Detection certainty measures
As described above, the adaptive path planner bases decisions on the certainty of the detection network to inspect uncertain detections that are possibly incorrect. This requires a correlation between the certainty and errors made by the detection network, since the certainty is used as probability whether an object is present or not. In this paper, we evaluate six different methods to calculate uncertainty. The first one is the standard detection certainty as returned by the detection network for the best scoring class, $c^{\textrm{yolo}}$. This value represents a combination of the certainty for the bounding box size and location, and the class of an object. However, most networks tend to over- or under-estimate the confidence (Gawlikowski et al., 2023). Therefore, we calculate five additional certainty measures using Monte-Carlo dropout (MCD). Calculating certainties using MCD is done in two steps: applying MCD on the YOLOv8 network (section 2.2.1), and calculating the certainty measures (section 2.2.2).

### v2

@@ Abstract
This paper presents an adaptive path planner for object search in agricultural fields using UAVs. The path planner uses a high-altitude coverage flight path and plans additional low-altitude inspections when the detection network is uncertain. The path planner was evaluated in an offline simulation environment containing real-world images. We trained a YOLOv8 detection network to detect artificial plants placed in grass fields to showcase the potential of our path planner. We evaluated the effect of different detection certainty measures, optimized the path planning parameters, investigated the effects of localization errors, and different numbers of objects in the field. The YOLOv8 detection confidence worked best to differentiate between true and false positive detections and was therefore used in the adaptive planner. The optimal parameters of the path planner depended on the distribution of objects in the field. When the objects were uniformly distributed, more low-altitude inspections were needed compared to a non-uniform distribution of objects, resulting in a longer path length. The adaptive planner proved to be robust against localization uncertainty. When increasing the number of objects, the flight path length increased, especially when the objects were uniformly distributed. When the objects were non-uniformly distributed, the adaptive path planner yielded a shorter path than a low-altitude coverage path, even with a high number of objects. Overall, the presented adaptive path planner allowed finding non-uniformly distributed objects in a field faster than a coverage path planner and resulted in a compatible detection accuracy. The path planner is made available at https://github.com/wur-abe/uav_adaptive_planner.
@@ 1 Introduction
Over the last years, the use of Unmanned Aerial Vehicles (UAVs) in agriculture has grown considerably (Rejeb et al., 2022). Applications are ranging from yield estimation (Yuan et al., 2024) to weed (Murad et al., 2023) and disease detection (Shahi et al., 2023; Kouadio et al., 2023). These applications require objects, such as weeds and diseases, to be detected in a large field. Typically, the UAV follows a pre-defined low-altitude coverage flight path at fixed altitude (Cabreira et al., 2019). These paths are easy to use and cover the whole field with an equal spatial resolution. However, for an object detection task, it is not needed to cover the whole field with an equal spatial resolution. Flying at a higher altitude will increase the field-of-view of the UAV and thereby decrease the total flight path length and associated flight time, but decrease the spatial resolution and thereby the detection accuracy. Only decreasing the altitude of the UAV when the detection is uncertain can decrease the flight time while achieving a similar detection as a low-altitude coverage path. To optimize the battery capacity of UAVs, it is important to use a method that minimizes the flight path length while maintaining a good detection accuracy. Especially when using UAVs for autonomous and frequent monitoring of fields, it is important to limit the flight time to increase the area that can be monitored using a single UAV.
@@
Adaptive path planning, sometimes called informative path planning, is a path planning approach that aims to plan a trajectory that maximizes gathering of task-relevant information while keeping within specific resource limits, such as battery capacity (Popović et al., 2024). This can be done by making the path planning reactive to real-time data, making sequential decisions about the UAV’s flight path based on newly gathered information (Popović et al., 2024). Adaptive path planning with UAVs is used, for instance, to detect anomalies in ocean topography (Blanchard and Sapsis, 2022), detection of weeds in agricultural fields (Popović et al., 2017), and multi-resolution semantic segmentation (Stache et al., 2023). Popović et al. (2017) showed that an adaptive path planner yielded a 50% lower map uncertainty (measured using entropy) in the same amount of time for active weed classification (simulated using printed markers) compared to a low-altitude coverage flight path. Stache et al. (2023) proposed a path planner for semantic segmentation that uses an estimated relation between altitude and segmentation accuracy. It performs close inspections at lower altitudes only when required and thereby minimizing the flight time. The path planner was demonstrated by segmenting soil, plants, and weeds by artificially flying over the WeedMap dataset and elements like asphalt, buildings, water, etc., by artificially flying over satellite images. They show a high segmentation accuracy while minimizing the flight time. van Essen et al. (2024) presents a learning based adaptive path planner that uses Reinforcement Learning to localize objects of interest in a field. The learned policy uses uncertain prior knowledge of the field together with object detections in the current field-of-view to control a UAV. In a simulated environment, they show that learned path outperforms a coverage planner by localizing 80% of the objects in less than half of the flight time. However, all these methods do not directly use the certainty estimation by the detection or segmentation network, even though such an estimate could be valuable for decision making in an adaptive path planner.
@@
When increasing the altitude of the UAV, the field-of-view increases, and thereby the spatial resolution of the image decreases. Due to a smaller number of pixels describing an object, the number of false positive and false negative detections will increase, and thereby the accuracy decreases. An adaptive path planner can exploit the trade-off between altitude and accuracy by only capturing low-altitude images when needed. For example, at a high altitude, the detection network may detect an object that appears to be a weed. Based on the uncertainty associated with the detection, a decision can be made to accept the detection as reliable or gather a new image at a lower altitude.
@@
A detection network typically outputs a confidence belonging to a detection as a measure of uncertainty. When the detection confidence of an object is low, an adaptive path planner can, for example, plan a low-altitude inspection of this object. However, neural networks tend to over- or under-estimate the confidence (Gawlikowski et al., 2023), leading to a rejection of correct detections and an inclusion of wrong detections. Several methods have been applied in literature to provide a better uncertainty estimation. Applying Monte-Carlo Dropout (MCD) during inference is often applied due to its low implementation effort (Gawlikowski et al., 2023). Myojin et al. (2019) showed a 1-2% increase in accuracy when using uncertainty measures estimated using MCD for detecting lunar craters on images using an adapted version of YOLOv3. In agriculture, MCD has been applied for uncertainty estimation in active learning for broccoli detection (Blok et al., 2022), fish species detection (Sokolova et al., 2024), obstacle detection in solar farms (Rodriguez-Vazquez et al., 2024), and assessing uncertainty in chicken plumage condition (Lamping et al., 2023). They illustrate that MCD can be used to estimate uncertainty in neural networks.
@@
The objective of this paper is to develop and evaluate an adaptive path planner that finds objects in agricultural fields faster than a low-altitude coverage path planner while maintaining a similar detection performance. This adaptive path planning method follows a high-altitude coverage flight path and uses the detection certainty to perform low-altitude inspections when the detection certainty is low. Compared to existing approaches in the literature, the developed adaptive path planner directly uses uncertainty estimates from an object detection network, making it more flexible for other applications. Besides the field boundaries, it does not need prior knowledge of the field, and the performance was evaluated on real-world data in combination with a real detection network. Furthermore, we open-sourced the code at https://github.com/wur-abe/uav_adaptive_planner. Specifically, we (1) examined the effect of different object detection certainty estimation methods calculated using MCD, (2) optimized the flight altitude and certainty thresholds of the planner, (3) analyzed the influence of errors in the UAV’s localization, and finally, (4) investigated the efficiency of the adaptive path planner to the number of objects in the field.
@@ 2.1 Adaptive path planner
The concept of the proposed adaptive path planner is visualized in Figure 1. First, the flight area to be covered by the UAV needs to be manually selected. Based on the flight area, a high-altitude coverage path is planned (section 2.1.1) and executed. At each waypoint, we take an image from a top-down perspective (nadir) and a detection network is used to detect the objects in real-time (section 2.1.2). The detected objects are then georeferenced (section 2.1.3) and mapped (section 2.1.4). After completing the high-altitude coverage path, an inspection path is planned with waypoints at a lower altitude at the locations of the object detections that were uncertain (section 2.1.5). To limit the number of waypoints for low-altitude inspection, the waypoints are filtered to remove overlapping waypoints. The inspection flight path is executed in a similar way as the high-altitude coverage path. During the inspection flight, the mapped objects are updated and previously unseen objects are added when needed. Finally, all object locations in the map (both from the high- and low-altitude flight) with a certainty larger than $c_{\textrm{eval}}$ are used for evaluation.
@@ 2.1.1 Planning of high-altitude coverage path
The first step of the proposed method is to manually select a flight area by drawing a polygon on a satellite image indicating the boundaries of the field. Within this area, a coverage flight path is planned using Fields2Cover (Mier et al., 2023) at an altitude of $h_{\textrm{cov}}$. The field-of-view, $\textrm{FOV}(h)$ (m) at altitude $h$ is calculated by:
$\textrm{FOV}(h)=\frac{\textbf{c}\odot h}{f},$ (1)
where $\odot$ denotes an element-wise multiplication, $\textbf{c}=[c_{w},c_{h}]^{T}$ the sensor width and height in mm, and $f$ the camera focal length in mm. However, when the localization of the UAV is imperfect, taking the calculated field-of-view as row width for the coverage path planner would then result in missing a part of the field. Therefore, to reduce the impact of localization errors, we set the row width to 90% of the field-of-view by adding an overlap of 5% of the field-of-view to both sides. Likewise, the distance between two waypoints on the flight path is set to 90% of the field-of-view. Figure 2 shows an example of the coverage flight paths for 12m, 24m, 36m, and 48m altitude.
Figure 2: Example of the coverage flight path for 12m (a), 24m (b), 36m (c), and 48m(d) altitude.
@@ 2.1.2 Detection
At each waypoint, an image is taken from a top-down perspective. In this image, objects are detected using YOLOv8. (Jocher et al., 2023). YOLOv8 is the eighth version of the YOLO (You Only Look Once) family of fully convolutional neural networks for object detection, which predicts both the bounding box coordinates as well as the class probabilities in a single pass through the network. The nano version of YOLOv8 is used, which was selected to have low hardware requirements, and allows for detection on board a UAV. During inference, an image input size of 2048x2048 pixels is used, and the minimum confidence is set to $c_{\textrm{reject}}$.
@@ 2.1.3 Georeferencing
Georeferencing object detections is done in real-time using the UAV’s position and heading. The first step is to calculate the field-of-view, $\textrm{FOV}(h)$, at the current altitude of the UAV using Equation 1. The pixel coordinates of the detected object $o$ in the image $[u_{o},v_{o}]^{T}$ (px) can be transformed in the local camera frame, $[x^{\textrm{local}}_{o},y^{\textrm{local}}_{o}]^{T}$ (m), by:
$\begin{bmatrix}x^{\textrm{local}}_{o}\\ y^{\textrm{local}}_{o}\end{bmatrix}=\Biggl{(}\begin{bmatrix}u_{o}\\ v_{o}\end{bmatrix}\odot\begin{bmatrix}\frac{1}{I_{w}}\\ \frac{1}{I_{h}}\end{bmatrix}-\frac{1}{2}\Biggr{)}\odot\textrm{FOV}(h),$ (2)
@@
where $\psi$ is the counter-clockwise heading of the drone relative with respect to the north and $[x^{\textrm{utm}}_{\textrm{uav}},y^{\textrm{utm}}_{\textrm{uav}}]^{T}$ the UAV’s UTM coordinates. Since UTM coordinates use the North-up orientation, we take the negative of $y^{\textrm{local}}_{o}$.
@@ 2.1.4 Mapping
The GNSS location of each detection is stored in a map during operation. Because of overlap between waypoints and flying at different altitudes, objects may be observed multiple times. To merge these observations, all detections with an Euclidean distance of less than $d_{\textrm{dist}}$ are considered the same object. The default value of $d_{\textrm{dist}}=0.35$m was selected based on the minimum distance between the objects in the field.
@@ 2.1.5 Planning of low-altitude inspection flight path
After executing the high-altitude coverage flight path, an inspection flight path is planned to observe objects in the map with a low certainty (see section 2.2) in more spatial detail at a lower altitude. For all the objects stored in the map, we make a decision $\mathcal{D}_{i}$ for each object $i$, to accept, reject, or have a closer inspection of the object using:
@@
where $c_{i}$ is the certainty corresponding to the object (see section 2.2), $c_{\textrm{accept}}$ the certainty threshold to accept an object without further inspection and $c_{\textrm{reject}}$ the certainty threshold to reject and ignore an object. An inspection flight path is planned at an altitude of $h_{\textrm{inspect}}$ to visit all objects where $\mathcal{D}_{i}$ equals ’inspect’. The resulting list of waypoints is ordered to get the shortest route using iterative stochastic local search by minimizing the total distance (Mulvad, 2022). For every waypoint, the heading is set towards the next waypoint. To shorten the created path, the waypoints are iteratively filtered by removing next waypoints when the corresponding detection is expected to be visible in the field-of-view of the current waypoint. Figure 3 shows an example of an inspection flight path.
@@ 2.2 Detection certainty measures
As described above, the adaptive path planner bases decisions on the certainty of the detection network to inspect uncertain detections that are possibly incorrect. This requires a correlation between the certainty and errors made by the detection network, since the certainty is used as a probability of whether an object is present or not. In this paper, we evaluate six different methods to calculate uncertainty. The first one is the standard detection certainty as returned by the detection network for the best scoring class, $c^{\textrm{yolo}}$. This value represents a combination of the certainty for the bounding box size and location, and the class of an object. However, most networks tend to over- or under-estimate the confidence (Gawlikowski et al., 2023). Therefore, we calculate five additional certainty measures using Monte-Carlo dropout (MCD). Calculating certainties using MCD is done in two steps: applying MCD on the YOLOv8 network (section 2.2.1), and calculating the certainty measures (section 2.2.2).